OpenStudy (anonymous):

solve the initial value problem tx''-x'=t^2, x (0)=0

OpenStudy (anonymous):

Notice \[ x_p(t)= \frac 1 3 t^3 \] is a particular solution of the HE

OpenStudy (anonymous):

Sorry is a particular solution of the NHE

OpenStudy (anonymous):

Find r so that \[ x(t) = t^r \] is a solution to HE

OpenStudy (anonymous):

You find r=0 and r =2 So the general equation of the HE is of the form \[ a + b t^2 \] So the solution is of the form \[ x(t)=a + b t^2+ \frac {t^3}3 \]

OpenStudy (anonymous):

since x(0) =0, then a=0; so \[ b t^2 +\frac {t^3}3 \]

OpenStudy (anonymous):

I meant \[ x(t)=b t^2 +\frac {t^3}3 \]

OpenStudy (anonymous):

You need another initial condition to find b

OpenStudy (anonymous):

How do i solve this using laplace transform?

OpenStudy (anonymous):

You're missing at least one other initial condition...

OpenStudy (anonymous):

@SithsAndGiggles , nope. that's the way the equation is set up.